# Flexible MOF Generation with Torsion-Aware Flow Matching

### v1

@@ Abstract
Designing metal–organic frameworks (MOFs) with novel chemistries is a long-standing challenge due to their large combinatorial space and the complex 3D arrangements of building blocks. While recent deep generative models have enabled scalable MOF generation, they assume (1) a fixed set of building blocks and (2) known ground-truth local block-wise 3D coordinates. However, this limits their ability to (1) design novel MOFs and (2) generate the structure using novel building blocks. We propose a two-stage de novo MOF generation framework that overcomes these limitations by modeling both chemical and geometric degrees of freedom. First, we train a SMILES-based autoregressive model to generate novel metal and organic building blocks, paired with cheminformatics for 3D structure initialization. Second, we introduce a flow-matching model that predicts translations, rotations, and torsional angles to assemble flexible blocks into valid 3D frameworks. Our experiments demonstrate improved reconstruction accuracy, the generation of valid, novel, and unique MOFs, and the ability of our model to create novel building blocks.
@@ 1 Introduction
Metal–organic frameworks (MOFs) are highly crystalline materials known for their permanent porosity, structural versatility, and broad applications in fields such as gas storage and separations [1, 2], catalysis [3], and drug delivery [4]. MOFs are designed by assembling chemical building blocks, i.e., metal clusters and organic linkers; their properties (e.g., pore size) can be tuned by swapping or modifying these components. This tunability has enabled the synthesis of over 100,000 distinct MOF structures to date [5], yet the theoretical space of possible MOFs is vastly larger, on the order of millions of structures [6]. Exploring this large design space is a grand challenge in materials science.
@@
[…]
Contribution. To address these limitations, we introduce MOFFlow-2, a two-stage generative framework capable of producing novel MOF chemistries and predicting high-fidelity 3D structures (Figure 1). Unlike prior methods, MOFFlow-2 does not rely on a fixed building block library or assume rigid conformations.
In the first stage, an autoregressive Transformer generates a canonicalized sequence of MOF building blocks represented as SMILES [13], enabling the creation of entirely novel chemistries. These SMILES can be initialized to 3D structures using cheminformatics tools such as RDKit [14]. In the second stage, a flow-based model predicts the full 3D MOF structure by jointly modeling (1) rotations and translations of building blocks, (2) lattice structure, and (3) torsion angle for rotatable bonds in the organic linkers. Crucially, by explicitly modeling the torsion angle, MOFFlow-2 captures the conformational flexibility of organic linkers and no longer depends on predefined and rigid 3D conformations. See Table 1 for comparison to prior works.
Figure 1: Overview of MOFFlow-2. MOFFlow-2 is a two-stage generative framework for MOF generation and structure prediction. The first stage uses a building block generator to generate an MOF sequence in SMILES representation, which is initialized to 3D coordinates with the metal library and RDKit. In the second stage, our structure prediction model assembles these building blocks by modeling translation $\bm{\tau}$, rotation $\bm{q}$, torsion $\bm{\phi}$, and lattice $\bm{\ell}$.
@@ 3.2 Flow matching on Riemannian manifolds
Flows on Riemannian manifolds. Here, we define CNF on Riemannian manifolds. Consider a smooth connected Riemannian manifold $\mathcal{M}$ with metric $g$, where each point $x\in\mathcal{M}$ has an associated tangent space $\mathcal{T}_{x}\mathcal{M}$ with inner product $\langle\cdot,\cdot\rangle_{g}$. A flow $\phi_{t}(x):\mathcal{M}\rightarrow{}\mathcal{M}$ is defined by the ordinary differential equation (ODE) $\frac{d}{dt}\phi_{t}(x)=u_{t}(\phi_{t}(x)),\phi_{0}(x)=x\sim p_{0}$ where $t\in[0,1]$ and $u_{t}(x)\in\mathcal{T}_{x}\mathcal{M}$ is the corresponding time-dependent vector field. We say that $u_{t}$ generates probability paths $p_{t}(x)$ if $p_{t}(x)=[\phi_{t}]_{\#}p_{0}(x)$, where ##\## denotes the pushforward operator.
@@
[…]
@@ 4.1 Autoregressive SMILES generation
[…]
Next, Given the canonical MOF sequence $\mathcal{B}_{\text{2D}}$, we tokenize it into a sequence of discrete symbols $[b_{1},\dots,b_{S}]$ by mapping substrings (e.g., atoms, bonds, or special tokens) into elements of a fixed vocabulary. We use a SMILES-aware regular expression [28] for tokenization, resulting in a vocabulary of size 59. See Section A.1 for an example of the tokenization process.
@@
Due to their complex chemistry, only a limited set of metal clusters is commonly used in MOFs [12]. Based on this observation, we construct a metal library of averaged 3D coordinates from the training data. Specifically, we group metal clusters by their canonical SMILES, select a random reference structure within each group, align all other structures to the reference using root mean square distance (RMSD) minimization, and compute the average of the aligned coordinates. [^1] Appendix B verifies this claim – our dataset contains only 7 distinct metal types with low RMSD variability.
@@ 5 MOF structure prediction with torsional degrees of freedom
Our MOF structure prediction model $p_{\theta}(\bm{q},\bm{\tau},\bm{\ell},\bm{\phi}|\mathcal{B}_{\text{3D}})$ is a flow-based model that predicts the 3D structural degrees of freedom – rotations $\bm{q}$, translations $\bm{\tau}$, torsion angles $\bm{\phi}$, and lattice parameters $\bm{\ell}$ – conditioned on the initialized 3D building block structure $\mathcal{B}_{\text{3D}}$. In this section, we present the training algorithm based on Riemannian flow matching, a Transformer-based architecture for predicting each structural component, and a canonicalization procedure for rotations and torsions to ensure consistent targets during training.
@@ 5.1 Training algorithm
Prior distribution. We define the priors over each structural component as follows. Rotations and torsions are sampled independently from uniform distributions on their respective manifolds: $q\sim\mathcal{U}(SO(3))$ and $\phi\sim\mathcal{U}(SO(2))$. Translations are drawn from a standard normal distribution with the center of mass removed: $\bm{\tau}\sim\mathcal{N}(0,I_{M})$. For the lattice parameters, the lengths $(a,b,c)$ are sampled independently from log-normal distributions: $(a,b,c)\sim\operatorname{LogNormal}(\bm{\mu},\bm{\sigma})$, where $\bm{\mu}=(\mu_{a},\mu_{b},\mu_{c})$ and $\bm{\sigma}=(\sigma_{a},\sigma_{b},\sigma_{c})$ are estimated by maximum likelihood objective; the lattice angles $(\alpha,\beta,\gamma)$ are sampled uniformly from the range $[60^{\circ},120^{\circ}]$ [9, 10].
Conditional flows, vector fields, and flow matching loss. Following Chen and Lipman [26], we define conditional flows on each manifold as geodesic interpolations between a prior sample $\bm{z}_{0}$ and a target data $\bm{z}_{1}$, where $\bm{z}_{1}$ represents one of translation, rotation, torsion, or lattice parameter. Specifically, we define the flow as $\bm{z}_{t}=\exp_{\bm{z}_{0}}(t\log_{\bm{z}_{0}}(\bm{z}_{1}))$, where $t\in[0,1]$ and $\log(\cdot)$, $\exp(\cdot)$ denote logarithm map and exponential map on respective manifolds (see Section C.1 manifold-specific details). This allows parameterizing the vector field by $v_{t}(\bm{z}_{t};\theta)={\log_{\bm{z}_{t}}(\hat{\bm{z}}_{1})}/(1-t)$ where the neural network predicts the clean data $\hat{\bm{z}}_{1}$ from intermediate state $\bm{z}_{t}$. Furthermore, the conditional vector field is given by $u_{t}(\bm{z}_{t}|\bm{z}_{1})={\log_{\bm{z}_{t}}(\bm{z}_{1})}/(1-t)$.
@@
where $\bm{X}_{t}$ is the noisy coordinate generated by applying rotation $\Delta\bm{q}_{1\rightarrow t}$, translation $\bm{\tau}_{t}$, and torsion $\bm{\phi}_{t}$ to the clean $\bm{X}_{1}$ (see Section C.2 for precise definition); $\Delta\bm{q}_{t\rightarrow 1}:=\bm{q}_{t}\bm{q}_{1}^{\top}$ is the rotation that aligns $\bm{X}_{t}$ back to $\bm{X}_{1}$; and the hat symbol $\hat{\ }$ indicates the model’s prediction of each variable. Under this formulation, we can rewrite the training objective as:
@@
where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ are scaling coefficients for rotation, translation, torsion, and lattice, respectively. We present detailed training and inference algorithms in Section C.3.
@@ 5.2 Model architecture
Our structure prediction model is built on a non-equivariant Transformer backbone. The architecture comprises an initialization module, an interaction module, and an output module (Figure 2)
Figure 2: Structure prediction model architecture. The model consists of three main modules: an initialization module that encodes atom features; an interaction module based on a Transformer encoder; and an output module with four prediction heads for each structural component – i.e., rotation (dimension $M$, the number of building blocks), translation ($M$), lattice parameters ($1$), and torsion angles (dimension $P$, the number of rotatable bonds). Notably, torsion angles are predicted by first constructing rotatable bond features with four corresponding dihedral atoms and updating the feature with attention to nearby atoms.
Initialization module. The initial node embeddings $h_{i}^{(0)}\in\mathbb{R}^{D}$ for $i$-th atom is computed as $\smash{h_{i}^{(0)}=[E_{\theta}(a_{i}),f(a_{i}),x_{i},\bm{\ell}_{t},\varphi(t)]+\varphi(k)}$, where $E_{\theta}(a_{i})$ is a learnable atom type embedding, $f(a_{i})$ is a binary feature vector encoding atom-level features such as aromaticity [15], and $\varphi(\cdot)$ is sinusoidal embedding function, $t$ is timestep, and $k$ is the index of the building block that contains the $i$-th atom. The building block index $k$ resolves ambiguity between blocks with identical SMILES by lexicographic ordering similar to Section 4.1. Full details are provided in Section D.1.
@@
To compute self-attention scores, we first compute the query, key, and value for each target atom $i$ and its neighbor atom $j$ as: $\mathsf{\mathbf{q}}_{i}\leftarrow{}\operatorname{Linear}_{Q}(h_{i})$, $\mathsf{\mathbf{k}}_{ij}\leftarrow{}\operatorname{Linear}_{K}([h_{j},e_{ij}])$, $\mathsf{\mathbf{v}}_{ij}\leftarrow{}\operatorname{Linear}_{V}([h_{j},e_{ij}])$. Then the attention score from $i$ to $j$ is computed as $a_{ij}=\operatorname{Softmax}_{j\in\mathcal{N}(i)}({\mathsf{\mathbf{q}}_{i}^{\top}\mathsf{\mathbf{k}}_{ij}}/{\sqrt{D}})$, where $D$ is the hidden dimension, and $\mathcal{N}(i)$ is the set of neighbors of atom $i$. Finally, we update the node feature by aggregating the value vectors from its neighbors: $h_{i}\leftarrow\sum_{j\in\mathcal{N}(i)}a_{ij}\mathsf{\mathbf{v}}_{ij}$. We provide additional architectural details and hyperparameters in Section D.2.
Output module. The output module consists of four heads that predict rotation $\Delta\hat{\bm{q}}_{1\rightarrow 1}$, translation $\hat{\bm{\tau}}_{1}$, torsion $\hat{\bm{\phi}}_{1}$, and lattice parameters ^$\hat{\bm{\ell}}_{1}$.
(Lattice parameter) To predict the lattice parameter, we mean-pool the node embeddings $\{h_{i}\}_{i=1}^{N}$ and apply a two-layer MLP with a GELU activation [30], i.e., set ^$\hat{\bm{\ell}}=\operatorname{MLP}(\frac{1}{N}\sum_{i=1}^{N}h_{i})$.
@@
(Torsion) To predict the torsion angles, we first construct invariant features for each rotatable bond and refine them with an attention mechanism. Specifically, for each rotatable bond between atoms $(j,k)$, we first define the dihedral angle using four atoms $(i,j,k,l)$, where atoms $i$ and $l$ are selected according to a consistent canonicalization scheme described in Section 5.3. The corresponding rotatable bond feature $h_{ijkl}$ is constructed by concatenating the atom features and applying a linear transformation in both forward and reverse directions:
@@ 5.3 Preprocessing with canonicalization and MOF matching
Canonicalization of rotation. MOF building blocks often have high symmetry point groups, where multiple rotations produce indistinguishable structures. As a result, there exist multiple “ground-truth” rotations $\Delta q_{t\rightarrow 1}$ that align $X_{t}$ to $X_{1}$, leading to unstable training (Figure 3(a)).
@@
To resolve this ambiguity, we introduce a canonicalization procedure that selects a unique rotation target for each symmetric building block. Given a clean conformation $\mathcal{C}_{1}=(A,X_{1})$ and a noised conformation $\mathcal{C}_{0}=(A,X_{0})$, we first identify the set of symmetry-preserving rotations of $\mathcal{C}_{0}$, defined as :=$\mathcal{R}:=\{\mathfrak{g}\in O(3)|\mathfrak{g}\cdot\mathcal{C}_{0}=\mathcal{C}_{0}\}\cap SO(3)$. We then apply each $\mathfrak{g}\in\mathcal{R}$ to $X_{0}$ to generate symmetrically equivalent conformations. Among these, we select the one with the lowest RMSD to $X_{1}$ and recompute the rotation between the two structures with Kabsch alignment [34] (Algorithm 1). This yields the canonical rotation $\Delta q_{1\rightarrow 0}^{*}$, which we interpolate to obtain the final target rotation $\Delta q_{t\rightarrow 1}$ for any $t\in[0,1]$.
Canonicalization of torsion. MOF building blocks frequently contain $\pi$-symmetry bonds associated with distinct rotations that results in an equivalent conformation [17, 35]. Therefore, representing torsion as a relative rotation can lead to multiple valid targets for a single noisy conformation. To address this, we adopt a torsion representation that explicitly specifies the four atoms of the dihedral, with the neighboring atoms selected consistently using RDKit’s canonical atom rankings [14]. Concretely, given a rotatable bond between atoms $j$ and $k$, we select atoms $i$ and $l$ as the lowest-ranked neighbors of $j$ and $k$, respectively (Figure 3(b)).
@@
MOF matching. We introduce a preprocessing step to reduce the distributional shift between the training and inference structures. Namely, during training, the model receives DFT-relaxed 3D conformations as input, while at inference, it receives coordinates initialized from the metal building block library and cheminformatics tools (see Section 4.2). These small discrepancies in bond lengths and angles leads to degraded performance [16, 15]. To mitigate this, we propose MOF matching, a preprocessing procedure adapted from conformer matching from Jing et al. [16]. Specifically, MOF matching replaces each metal building block with its template structure from the library and each organic linker with an RDKit-generated structure whose torsion angles are optimized to closely match the original conformation. The resulting matched coordinates are used for training to ensure consistency with inference-time input. See Appendix E for details.

### v3

@@ Abstract
Designing metal-organic frameworks (MOFs) with novel chemistries is a longstanding challenge due to their large combinatorial space and complex 3D arrangements of the building blocks. While recent deep generative models have enabled scalable MOF generation, they assume (1) a fixed set of building blocks and (2) known local 3D coordinates of building blocks. However, this limits their ability to (1) design novel MOFs and (2) generate the structure using novel building blocks. We propose a two-stage MOF generation framework that overcomes these limitations by modeling both chemical and geometric degrees of freedom. First, we train an SMILES-based autoregressive model to generate metal and organic building blocks, paired with a cheminformatics toolkit for 3D structure initialization. Second, we introduce a flow matching model that predicts translations, rotations, and torsional angles to assemble the blocks into valid 3D frameworks. Our experiments demonstrate improved reconstruction accuracy, the generation of valid, novel, and unique MOFs, and the ability to create novel building blocks. Our code is available at https://github.com/nayoung10/MOFFlow-2.
@@ 1 Introduction
Metal-organic frameworks (MOFs) are highly crystalline materials known for their permanent porosity, structural versatility, and broad applications in fields such as gas storage and separations [1, 2], catalysis [3], and drug delivery [4]. MOFs are designed by assembling chemical building blocks, i.e., metal clusters and organic linkers; their properties (e.g., pore size) can be tuned by swapping or modifying these components. This tunability has enabled the synthesis of over 100,000 distinct MOF structures to date [5], yet the theoretical space of possible MOFs is vastly larger, on the order of millions of structures [6]. Exploring this large design space is a grand challenge in materials science.
@@
MOFDiff [11]
MOFFlow [10] -
MOFFlow-2
Contribution. To address these limitations, we introduce MOFFlow-2, a two-stage generative framework capable of producing novel MOF chemistries and predicting high-fidelity 3D structures (Figure˜1). Unlike prior methods, MOFFlow-2 does not rely on a fixed building block library or assume rigid conformations.
In the first stage, an autoregressive Transformer generates a canonicalized sequence of MOF building blocks represented as SMILES [13], enabling the creation of entirely novel chemistries. These SMILES can be initialized to 3D structures using cheminformatics tools such as RDKit [14]. In the second stage, a flow-based model predicts the full 3D MOF structure by jointly modeling (1) rotations and translations of building blocks, (2) lattice structure, and (3) torsion angle for rotatable bonds in the organic linkers. Crucially, by explicitly modeling the torsion angle, MOFFlow-2 captures the conformational flexibility of organic linkers and no longer depends on predefined and rigid 3D conformations. See Table˜1 for comparison to prior works.
@@ 3.2 Flow matching on Riemannian manifolds
Flows on Riemannian manifolds. Here, we define CNF on Riemannian manifolds. Consider a smooth connected Riemannian manifold $\mathcal{M}$ with metric $g$, where each point $x\in\mathcal{M}$ has an associated tangent space $\mathcal{T}_{x}\mathcal{M}$ with inner product $\langle\cdot,\cdot\rangle_{g}$. A flow $\phi_{t}(x):\mathcal{M}\rightarrow{}\mathcal{M}$ is defined by the ordinary differential equation (ODE) $\frac{d}{dt}\phi_{t}(x)=u_{t}(\phi_{t}(x)),\phi_{0}(x)=x\sim p_{0}$ where $t\in[0,1]$ and $u_{t}(x)\in\mathcal{T}_{x}\mathcal{M}$ is the corresponding time-dependent vector field. We say that $u_{t}$ generates probability paths $p_{t}(x)$ if $p_{t}(x)=[\phi_{t}]_{\#}p_{0}(x)$, where #\# denotes the pushforward operator.
@@
$p_{t}(x)=\int_{\mathcal{M}}p_{t}(x|x_{1})q(x_{1})\,d\mathrm{vol}(x_{1}),\quad u_{t}(x)=\int_{\mathcal{M}}u_{t}(x|x_{1})\frac{p_{t}(x|x_{1})q(x_{1})}{p_{t}(x)}\,d\mathrm{vol}(x_{1}),$ (2)
@@ 4.1 Autoregressive SMILES generation
Defining MOF sequence. Given an MOF with $M_{1}$ metal clusters $m_{1},\dots,m_{M_{1}}$ and $M_{2}$ organic linkers $o_{1},\dots,o_{M_{2}}$, we impose a canonical ordering on the building blocks to ensure a consistent sequence representation. Specifically, (1) Metal building blocks precede organic ones, separated by a special token <SEP>, (2) Within each group (metal or organic), building blocks are sorted by ascending molecular weight and separated with the symbol ".", and (3) The full sequence is wrapped with start and end tokens <BOS> and <EOS>. The resulting sequence is ${\mathcal{B}_{\text{2D}}}=\bigl[\texttt{<BOS>}~\texttt{m\_1.m\_2.}\cdots\texttt{.m\_M1}~\texttt{<SEP>}~\texttt{o\_1.o\_2.}\cdots\texttt{.o\_{M2}}~\texttt{<EOS>}\bigr]$, where m_i and o_j denote the SMILES string of metal and organic building blocks, respectively.
Next, Given the canonical MOF sequence $\mathcal{B}_{\text{2D}}$, we tokenize it into a sequence of discrete symbols $b_{1},\dots,b_{S}]$ by mapping substrings (e.g., atoms, bonds, or special tokens) into elements of a fixed vocabulary. We use a SMILES-aware regular expression [28] for tokenization, resulting in a vocabulary of size 59. See Section˜A.1 for an example of the tokenization process.
@@
Due to their complex chemistry, only a limited set of metal clusters is commonly used in MOFs [12]. Based on this observation, we construct a metal library of averaged 3D coordinates from the training data. Specifically, we group metal clusters by their canonical SMILES, select a random reference structure within each group, align all other structures to the reference using root mean square distance (RMSD) minimization, and compute the average of the aligned coordinates. [^1] Appendix B confirms this – our dataset has only 7–8 metal types with low RMSD variability.
@@ 5 MOF structure prediction with torsional degrees of freedom
[…]
@@ 5.1 Training algorithm
Prior distribution. We define the priors over each structural component as follows. Rotations and torsions are sampled independently from uniform distributions on their respective manifolds: $q\sim\mathcal{U}(SO(3))$ and $\phi\sim\mathcal{U}(SO(2))$. Translations are drawn from a standard normal distribution with the center of mass removed: $\bm{\tau}\sim\mathcal{N}(0,I_{M})$. For the lattice parameters, the lengths $(a,b,c)$ are sampled independently from log-normal distributions: $(a,b,c)\sim\operatorname{LogNormal}(\bm{\mu},\bm{\sigma})$, where $\bm{\mu}=(\mu_{a},\mu_{b},\mu_{c})$ and $\bm{\sigma}=(\sigma_{a},\sigma_{b},\sigma_{c})$ are estimated by maximum likelihood objective; the lattice angles $(\alpha,\beta,\gamma)$ are sampled uniformly from the range $60^{\circ},120^{\circ}] [$9, 10].
Conditional flows, vector fields, and flow matching loss. Following Chen and Lipman [26], we define conditional flows on each manifold as geodesic interpolations between a prior sample $\bm{z}_{0}$ and a target data $\bm{z}_{1}$, where $\bm{z}_{1}$ represents one of translation, rotation, torsion, or lattice parameter. Specifically, we define the flow as $\bm{z}_{t}=\exp_{\bm{z}_{0}}(t\log_{\bm{z}_{0}}(\bm{z}_{1}))$, where $t\in[0,1]$ and $\log(\cdot)$, $\exp(\cdot)$ denote logarithm map and exponential map on respective manifolds (see Section˜C.1 manifold-specific details). This allows parameterizing the vector field by $v_{t}(\bm{z}_{t};\theta)={\log_{\bm{z}_{t}}(\hat{\bm{z}}_{1})}/(1-t)$ where the neural network predicts the clean data $\hat{\bm{z}}_{1}$ from intermediate state $\bm{z}_{t}$. Furthermore, the conditional vector field is given by $u_{t}(\bm{z}_{t}|\bm{z}_{1})={\log_{\bm{z}_{t}}(\bm{z}_{1})}/(1-t)$.
@@
where $\bm{X}_{t}$ is the noisy coordinate generated by applying rotation $\Delta\bm{q}_{1\rightarrow t}$, translation $\bm{\tau}_{t}$, and torsion $\bm{\phi}_{t}$ to the clean $\bm{X}_{1}$ (see Section˜C.2 for precise definition); $\Delta\bm{q}_{t\rightarrow 1}:=\bm{q}_{t}\bm{q}_{1}^{\top}$ is the rotation that aligns $\bm{X}_{t}$ back to $\bm{X}_{1}$; and the hat symbol $\hat{\ }$ indicates the model’s prediction of each variable. Under this formulation, we can rewrite the training objective as:
@@
where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ are scaling coefficients for rotation, translation, torsion, and lattice, respectively. We present detailed training and inference algorithms in Section˜C.3.
@@ 5.2 Model architecture
Our structure prediction model is built on a non-equivariant Transformer backbone. The architecture comprises an initialization module, an interaction module, and an output module (Figure˜2)
Figure 2: Structure prediction model architecture. The model consists of three main modules: an initialization module that encodes atom features; an interaction module based on a Transformer encoder; and an output module with four prediction heads for each structural component – i.e., rotation (dimension $M$, the number of building blocks), translation ($M$), lattice parameters (11), and torsion angles (dimension $P$, the number of rotatable bonds). Notably, torsion angles are predicted by first constructing rotatable bond features with four corresponding dihedral atoms and updating the feature with attention to nearby atoms.
Initialization module. The initial node embeddings $h_{i}^{(0)}\in\mathbb{R}^{D}$ for $i$-th atom is computed as $\smash{h_{i}^{(0)}=[E_{\theta}(a_{i}),f(a_{i}),x_{i},\bm{\ell}_{t},\varphi(t)]+\varphi(k)}$, where $E_{\theta}(a_{i})$ is a learnable atom type embedding, $f(a_{i})$ is a binary feature vector encoding atom-level features such as aromaticity [15], and $\varphi(\cdot)$ is sinusoidal embedding function, $t$ is timestep, and $k$ is the index of the building block that contains the $i$-th atom. The building block index $k$ resolves ambiguity between blocks with identical SMILES by lexicographic ordering similar to Section˜4.1. Full details are provided in Section˜D.1.
@@
To compute self-attention scores, we first compute the query, key, and value for each target atom $i$ and its neighbor atom $j$ as: $\mathsf{\mathbf{q}}_{i}\leftarrow{}\operatorname{Linear}_{Q}(h_{i})$, $\mathsf{\mathbf{k}}_{ij}\leftarrow{}\operatorname{Linear}_{K}([h_{j},e_{ij}])$, $\mathsf{\mathbf{v}}_{ij}\leftarrow{}\operatorname{Linear}_{V}([h_{j},e_{ij}])$. Then the attention score from $i$ to $j$ is computed as $a_{ij}=\operatorname{Softmax}_{j\in\mathcal{N}(i)}({\mathsf{\mathbf{q}}_{i}^{\top}\mathsf{\mathbf{k}}_{ij}}/{\sqrt{D}})$, where $D$ is the hidden dimension, and $\mathcal{N}(i)$ is the set of neighbors of atom $i$. Finally, we update the node feature by aggregating the value vectors from its neighbors: $h_{i}\leftarrow\sum_{j\in\mathcal{N}(i)}a_{ij}\mathsf{\mathbf{v}}_{ij}$. We provide additional architectural details and hyperparameters in Section˜D.2.
Output module. The output module consists of four heads that predict rotation $\Delta\hat{\bm{q}}_{1\rightarrow 1}$, translation $\hat{\bm{\tau}}_{1}$, torsion $\hat{\bm{\phi}}_{1}$, and lattice parameters $\hat{\bm{\ell}}_{1}$.
(Lattice parameter) To predict the lattice parameter, we mean-pool the node embeddings $\{h_{i}\}_{i=1}^{N}$ and apply a two-layer MLP with a GELU activation [30], i.e., set $\hat{\bm{\ell}}=\operatorname{MLP}(\frac{1}{N}\sum_{i=1}^{N}h_{i})$.
@@
(Torsion) To predict the torsion angles, we first construct invariant features for each rotatable bond and refine them with an attention mechanism. Specifically, for each rotatable bond between atoms $(j,k)$, we first define the dihedral angle using four atoms $(i,j,k,l)$, where atoms $i$ and $l$ are selected according to a consistent canonicalization scheme described in Section˜5.3. The corresponding rotatable bond feature $h_{ijkl}$ is constructed by concatenating the atom features and applying a linear transformation in both forward and reverse directions:
@@ 5.3 Preprocessing with canonicalization and MOF matching
Canonicalization of rotation. MOF building blocks often have high symmetry point groups, where multiple rotations produce indistinguishable structures. As a result, there exist multiple “ground-truth” rotations $\Delta q_{t\rightarrow 1}$ that align $X_{t}$ to $X_{1}$, leading to unstable training (Figure˜3(a)).
@@
To resolve this ambiguity, we introduce a canonicalization procedure that selects a unique rotation target for each symmetric building block. Given a clean conformation $\mathcal{C}_{1}=(A,X_{1})$ and a noised conformation $\mathcal{C}_{0}=(A,X_{0})$, we first identify the set of symmetry-preserving rotations of $\mathcal{C}_{0}$, defined as $\mathcal{R}:=\{\mathfrak{g}\in O(3)|\mathfrak{g}\cdot\mathcal{C}_{0}=\mathcal{C}_{0}\}\cap SO(3)$. We then apply each $\mathfrak{g}\in\mathcal{R}$ to $X_{0}$ to generate symmetrically equivalent conformations. Among these, we select the one with the lowest RMSD to $X_{1}$ and recompute the rotation between the two structures with Kabsch alignment [34] (Algorithm˜1). This yields the canonical rotation $\Delta q_{1\rightarrow 0}^{*}$, which we interpolate to obtain the final target rotation $\Delta q_{t\rightarrow 1}$ for any $t\in[0,1]$.
Canonicalization of torsion. MOF building blocks frequently contain $\pi$-symmetry bonds associated with distinct rotations that results in an equivalent conformation [17, 35]. Therefore, representing torsion as a relative rotation can lead to multiple valid targets for a single noisy conformation. To address this, we adopt a torsion representation that explicitly specifies the four atoms of the dihedral, with the neighboring atoms selected consistently using RDKit’s canonical atom rankings [14]. Concretely, given a rotatable bond between atoms $j$ and $k$, we select atoms $i$ and $l$ as the lowest-ranked neighbors of $j$ and $k$, respectively (Figure˜3(b)).
@@
MOF matching. We introduce a preprocessing step to reduce the distributional shift between the training and inference structures. Namely, during training, the model receives DFT-relaxed 3D conformations as input, while at inference, it receives coordinates initialized from the metal building block library and cheminformatics tools (see Section˜4.2). These small discrepancies in bond lengths and angles leads to degraded performance [16, 15]. To mitigate this, we propose MOF matching, a preprocessing procedure adapted from conformer matching from Jing et al. [16]. Specifically, MOF matching replaces each metal building block with its template structure from the library and each organic linker with an RDKit-generated structure whose torsion angles are optimized to closely match the original conformation. The resulting matched coordinates are used for training to ensure consistency with inference-time input. See Appendix˜E for details.